OpenStudy (anonymous):

13. Model the function rule y = -2x - 3 with a table of values. Show your work. Can some one please help me out?

OpenStudy (anonymous):

Well, just pick a bunch of values for 'x' and evaluate the function for it. for example x=5 means y = -2*5 - 3 = -10 - 3 = -13 Then for finding the y intercept just place x = 0. For x intercept place y = 0 (and find the value of 'x') Can do?

OpenStudy (anonymous):

i have to make a table too dont know how to do none of that can you give me a explain

OpenStudy (anonymous):

if i have an explain how to do i can do it

OpenStudy (anonymous):

Well, there is an identical post already so: http://openstudy.com/study#/updates/4ffdcc88e4b09082c06f1da4

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you use any number you what or do you have to use 0 1 2

OpenStudy (anonymous):

numbers around 0 are common for such problems. but you can pick others as well =] I suggest in addition and not instead though.